# FlippedRAG: Black-Box Opinion Manipulation Adversarial Attacks to Retrieval-Augmented Generation Models

### v3

@@ Abstract.
Retrieval-Augmented Generation (RAG) enriches LLMs by dynamically retrieving external knowledge, reducing hallucinations and satisfying real-time information needs. While existing research mainly targets RAG’s performance and efficiency, emerging studies highlight critical security concerns. Yet, current adversarial approaches remain limited, mostly addressing white-box scenarios or heuristic black-box attacks without fully investigating vulnerabilities in the retrieval phase. Additionally, prior works mainly focus on factoid Q&A tasks, their attacks lack complexity and can be easily corrected by advanced LLMs. In this paper, we investigate a more realistic and critical threat scenario: adversarial attacks intended for opinion manipulation against black-box RAG models, particularly on controversial topics. Specifically, we propose FlippedRAG, a transfer-based adversarial attack against black-box RAG systems. We first demonstrate that the underlying retriever of a black-box RAG system can be reverse-engineered and approximated by enumerating critical queries, candidates, and answers, enabling us to train a surrogate retriever. Leveraging the surrogate retriever, we further craft target poisoning triggers, altering vary few documents to effectively manipulate both retrieval and subsequent generation, transferring the attack to the original black-box RAG model. Extensive empirical results show that FlippedRAG substantially outperforms baseline methods, improving the average attack success rate by 16.7%. Across four diverse domains, FlippedRAG achieves on average a 50% directional shift in the opinion polarity of RAG-generated responses, ultimately causing a notable 20% shift in user cognition. Furthermore, we actively evaluate the performance of several potential defensive measures, concluding that existing mitigation strategies remain insufficient against such sophisticated manipulation attacks. These results highlight an urgent need for developing innovative defensive solutions to ensure the security and trustworthiness of RAG systems.
RAG, Black-box Adversarial Attack, Opinion Manipulation
*Corresponding author
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application securityccs: Information systems Information retrievalccs: Computing methodologies Natural language processing
@@ 1. Introduction
Retrieval-Augmented Generation (RAG) combines information retrieval with the generative capabilities of LLMs, enhancing the timeliness of knowledge acquisition and effectively mitigating the hallucination problem of these models (Gao et al., 2023; Zhao et al., 2024b).
@@
As the application scope of RAG expands, concerns about its security are also increasing. The basic RAG process typically consists of three components: a knowledge base (corpus collection), a retriever, and an LLM. Vulnerabilities arise when some of the retrieved documents are compromised, or prompt are tampered by malicious attackers, a scenario referred to as a manipulation attack against RAG. Previous studies have explored various forms of adversarial manipulation attacks, such as adversarial attack on the retriever (Liu et al., 2023b, c), prompt injection attack (Cai et al., 2022; Liu et al., 2023a; Jain et al., 2023), jailbreak attack for LLM (Deng et al., 2023; Li et al., 2023; Zhao et al., 2024a), and poisoning attack targeting the retrieval corpus in RAG (Cho et al., 2024; Zou et al., 2024; Xue et al., 2024).
Figure 1. (a) ChatGPT can be induced through user prompts to fully replicate the context segments referenced in its responses. Based on this observation, we can transparentize the retriever of the black-box RAG model by enumerating critical queries and candidates. (b) An overview of our black-box opinion manipulation adversarial attacks to RAG models and its potential impact on user’s cognition.
@@
However, previous studies (Zou et al., 2024; Xue et al., 2024; Cho et al., 2024) have primarily focused on attacks against RAG systems, which lack feasibility and have fundamental limitations. They mostly address white-box scenario or employ heuristic-based black-box attacks, without thoroughly analyzing the vulnerabilities within the retrieval stage. Furthermore, earlier research has mainly concentrated on factoid question answering tasks, such as questions like “Who is the CEO of OpenAI?” Those factual queries are less impactful, as RAG systems combined with fact-checking and value alignment mechanisms in LLMs are capable of rectifying these errors (Xiang et al., 2024; Tu et al., 2025).
@@
Differently, the scenario explored in this paper primarily addresses controversial and opinion-based inquiries in RAG, such as “Should abortion be legal?” These questions are open-ended and require advanced levels of logical reasoning and summarization capabilities from LLMs. Current research on controversial topics is limited. In scenarios involving controversial topics, opinion manipulation can create “Information Bubbles”, which subsequently lead to the homogenization of user opinions (Yue et al., 2023). In such cases, individuals’ views become easily swayed by the stance of the information they are exposed to. Attacks that manipulate opinions on such questions could potentially inflict deeper harm (Epstein and Robertson, 2015). Through carefully crafted inputs, attackers can influence the orientation of content generated by RAG models, thus jeopardizing users’ cognitive processes and decision-making abilities.
@@
Thus, this paper primarily investigates adversarial opinion manipulation targeting the retriever in black-box RAG systems, which align with realistic and practical scenarios. The threat model presented here can be characterized as follows: the adversary can only query the RAG system and does not have access to the complete knowledge base or corpus, the retriever, or the parameters of the RAG. The attacker is only capable of injecting limited adversarially modified candidate texts into the corpus, while the retriever and the LLM remain black-boxed, intact, and unmodifiable.
@@
To address the aforementioned challenges, in this paper, we propose FlippedRAG, a black-box attack method, to explore the reliability of RAG in controversial topics and investigate its impact on user cognition, as shown in Figure 1. By employing black-box retriever imitation (Liu et al., 2023b) for opinion manipulation, we reveal specific vulnerabilities in RAG systems. Our motivation is based on the observation in Figure 1(a). When we queried the RAG on a specific topic and included in the user prompt an instruction requesting LLMs to output the original content of its referenced retrieval results, it would faithfully reproduce the exact text segments it referenced. The reproduced outputs precisely matched the corresponding content on the web pages linked in the reference citations and reflect the functionality of relevance matching. It allows us to design a retriever imitation method for black-box RAG to address the limitations of white-box attacks or heuristic-based attacks against such systems. Specifically, we utilize carefully designed instructions to obtain the retrieval results within the RAG model. Subsequently, based on the constructed pairs (query, candidate) from the retrieval list, we train a surrogate model using the extracted retrieval data (Liu et al., 2023b; Wu et al., 2022b) to approximate and transparentize the relevance preferences of the retriever in the black-box RAG system.
@@
Based on this surrogate model, we develop an attack strategy aimed at manipulating the opinions of candidate documents. By attacking this surrogate model, we generate adversarial opinion manipulation triggers, which are then transferred to the target RAG corpus, as shown in Figure 1 (b). We conduct experiments on opinion datasets encompassing multiple topics to validate the effectiveness and scope of the attack strategy without relying on internal knowledge of the RAG model. The experimental results reveal that the proposed attack strategy can significantly alter the opinion polarity of the content generated by RAG. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack. This not only highlights the vulnerability of the retrieval model and RAG model but, more importantly, underscores the potential negative impact on user cognition and decision-making. Such vulnerabilities can increase the risk of misleading users into accepting incorrect or biased information.
@@
(1) We propose FlippedRAG, a novel transfer-based opinion manipulation attack against black-box RAG systems. It realistically reflects practical threat scenarios, and to our knowledge, we are among the first to study adversarial opinion manipulation targeting open-ended and controversial topics in RAG-based models.
(2) We propose a new approach to extract context documents and constructing pseudo-relevant training data for surrogate retriever models in black-box settings. It highlights risks regarding context data leakage in LLM applications. Extensive automated evaluations combined with user studies confirm that our method effectively manipulates the opinions generated by RAG models, ultimately influencing user perception toward specifically targeted viewpoints.
@@ 2. Related Works
2.1. Reliability of Neural Network Models
Research on the reliability of neural network models has been extensively studied. In 2013, Szegedy et al. (Szegedy et al., 2013) found that applying imperceptible perturbations to a neural network model during a classification task was sufficient to cause classification errors in CV. Later, scholars observed similar phenomena in NLP. Robin et al. (Jia and Liang, 2017) found that inserting perturbed text into original paragraphs significantly distracts computer systems without altering the correct answer or misleading humans. It reflects the robustness of neural network models, i.e., the ability to output stable and correct predictions in tackling the imperceptible additive noises (Wang et al., 2019). For LLMs, Wang et al. (Wang et al., 2023) proposed a comprehensive trustworthiness evaluation framework, assessing their reliability from various perspectives such as toxicity, adversarial robustness, stereotype bias, and fairness. While LLMs have greater capabilities compared with general deep neural network models, they also raise more concerns regarding security and reliability.
@@ 2.2. Adversarial Retrieval Attack Strategy
The adversarial retrieval attack strategy starts with manipulation at the word level. Under white-box setting, Ebrahimi et al. (Ebrahimi et al., 2017) utilize an atomic flip operation, which swaps one token for another, to generate adversarial text, named Hotflip. Hotflip eliminates reliance on rules, but the adversarial trigger it generates usually has incomplete semantics and insufficient grammar fluency. While it can deceive the target model, it cannot evade perplexity-based defenses. Wu et al. (Wu et al., 2022b) also proposed a word substitution ranking attack method called PRADA. To enhance the readability and effectiveness of the trigger, Song et al. (Song et al., 2020) propose an adversarial method under white-box setting, named Collision, which uses gradient optimization and beam search to produce the adversarial trigger. Collision method further imposes a soft constraint on collision generation by integrating a language model, reducing the perplexity of the collision. Futhermore, Liu et al. (Liu et al., 2023b) propose the Pairwise Anchor-based Trigger (PAT) generation under black-box setting. By adding the fluency and semantic consistency constraints, it generates adversarial triggers by optimizing the pairwise loss of top candidates and target candidates with triggers.
@@ 2.3. Reliability of RAG
Some studies have investigated the adversarial poisoning attacks on RAG models, including PoisonedRAG(Zou et al., 2024), BadRAG(Xue et al., 2024), Phantom(Chaudhari et al., 2024), and GARAG(Cho et al., 2024). The distinctive characteristics of these attack methodologies are summarized in the Table 1. “LLM Capability Utilization” refers to the motivation of leveraging LLMs’ capabilities to induce malicious outputs. “Depth” denotes the cognitive penetration depth of RAG attacks. Although both BadRAG(Xue et al., 2024) and Phantom(Chaudhari et al., 2024) target user perception, their manipulations remain confined to sentiment level without achieving deeper ideological persuasion.
As RAG is designed to overcome the hallucination problem in LLMs and enhance their generative capabilities, the reliability of content generated by RAG also becomes a major concern. Zhang et al. (Zhang et al., 2024) sought to identify the weaknesses of RAG by analyzing critical components in order to facilitate the injection of the attack sequence and crafting the malicious documents with a gradient-guided token mutation technique. Xiang et al. (Xiang et al., 2024) proposed an isolate-then-aggregate strategy, named RobustRAG, which first obtains responses from each passage of the LLM in isolation and then securely aggregates these isolated responses, to construct the first defense framework against retrieval corruption attacks. These studies are based on white-box scenarios and primarily focus on the robustness of RAG against corrupted and toxic content.
Table 1. Comparison of existing RAG attacks. DoS denotes denial-of-service.
@@
PoisonedRAG (Zou et al., 2024) Closed-ended, Factoid White-box/Black-box Logical reasoning
BadRAG (Xue et al., 2024) Open-ended, Factoid, DoS White-box Logical reasoning Sentiment
Phantom (Chaudhari et al., 2024) Open-ended, Factoid, DoS White-box Instruction-following Sentiment
GARAG (Cho et al., 2024) Closed-ended, Factoid Gray-box Noise resistence
FlippedRAG Controversial, Open-ended Black-box Induction, logical reasoning Perspectives
This paper proposes the use of adversarial retrieval attack to manipulate the retrieval process in black-box RAG, ensuring that corrupted documents with a specific opinion are ranked as highly as possible, thereby inducing the LLM responses to align with that opinion on controversial topics.
@@ 3.1. Overview
Our objective is to manipulate the opinions expressed in responses generated by black-box RAG models on controversial topics, with a focus on both the retrieval model and the LLM that performs the integrated generation task. Zhang et al. (Zhang et al., 2024) attempted to poison context documents to mislead the LLM into generating incorrect content. However, this approach necessitates extensive internal details of the LLM application, rendering it less feasible in real-world scenarios. In the black-box RAG context, the attacker lacks access to internal information of the RAG, including model architecture and scoring functions, and can only interact with the inputs and outputs of the RAG. Specifically, the attacker can only use the interface of the RAG and not directly access the retriever. Since the inputs to LLMs are comprised of the user query and context documents, and the user query is immutable, our attack strategy focuses on modifying candidate documents within the corpus. Although the attacker does not have access to the entire corpus, they can insert adversarially modified candidate texts into it, as many RAG applications source information from the Internet, where the content is publicly editable.
The basic framework of RAG comprises two components: the retriever and the LLM. These modules are serially connected, with the retriever sourcing context information from a knowledge base, upon which the LLM then performs the generation task. In black-box scenarios, attackers cannot modify the system prompts of the generative LLM, making it challenging to directly influence the generation results by exploiting any reliability flaws within the LLM itself. Therefore, we focus on exploiting the reliability flaws of the retriever to manipulate the retrieval ranking results. We employ specific instructions to induce RAG to reveal the context information it references. With this information, we train a surrogate model that approximates the black-box retriever, effectively transparentize it into a white-box model. We then employ this white-box model to generate adversarial triggers. By adding these adversarial triggers to candidate documents that reflect the target opinion, we enhance their relevance to the user query, increasing the likelihood that they will be included in the context passed to the generative LLM. Leveraging the strong capability of LLM for context understanding, summarization and instructions-following, we guide it to generate responses that align with the expected opinion. An overview of this process is shown in Figure 2. The proposed method is outlined in Algorithm 1 in Appendix A.1.
Figure 2. The overview of FlippedRAG for manipulating the opinions of RAG-generated content under black-box setting.
[…]
FlippedRAG Attack Method Overview.

### v4

@@ Abstract.
Retrieval-Augmented Generation (RAG) enriches LLMs by dynamically retrieving external knowledge, reducing hallucinations and satisfying real-time information needs. While existing research mainly targets RAG’s performance and efficiency, emerging studies highlight critical security concerns. Yet, current adversarial approaches remain limited, mostly addressing white-box scenarios or heuristic black-box attacks without fully investigating vulnerabilities in the retrieval phase. Additionally, prior works mainly focus on factoid Q&A tasks, their attacks lack complexity and can be easily corrected by advanced LLMs. In this paper, we investigate a more realistic and critical threat scenario: adversarial attacks intended for opinion manipulation against black-box RAG models, particularly on controversial topics. Specifically, we propose FlippedRAG, a transfer-based adversarial attack against black-box RAG-like systems. We first demonstrate that the underlying retriever of a black-box RAG can be reverse-engineered and approximated by enumerating critical queries, candidates, and answers, enabling us to train a surrogate retriever. Leveraging the surrogate retriever, we further craft target poisoning triggers, altering vary few documents to effectively manipulate both retrieval and subsequent generation, transferring the attack to the original black-box RAG model. Extensive empirical results show that FlippedRAG substantially outperforms baseline methods, improving the average attack success rate by 16.7%. Across four diverse domains, FlippedRAG achieves on average a 50% directional shift in the opinion polarity of RAG-generated responses, ultimately causing a notable 20% shift in user cognition. Furthermore, we actively evaluate the performance of several potential defensive measures, concluding that existing mitigation strategies remain insufficient against such sophisticated manipulation attacks. These results highlight an urgent need for developing innovative defensive solutions to ensure the security and trustworthiness of RAG systems.
RAG; Black-box Adversarial Attack; Opinion Manipulation
$c$s: Security and privacy Domain-specific security and privacy architecturesccs: Information systems Information extractionccs: Computing methodologies Natural language processing
@@ 1. Introduction
Retrieval-Augmented Generation (RAG) combines information retrieval with the generative capabilities of LLMs, enhancing the timeliness of knowledge acquisition and effectively mitigating the hallucination problem of these models (gao2023retrieval; zhao2024retrieval). As the application scope of RAG expands, concerns about its security are also increasing. The basic RAG process typically consists of three components: a knowledge base (corpus collection), a retriever, and an LLM. Vulnerabilities arise when some of the retrieved documents are compromised, or prompt are tampered by malicious attackers, a scenario referred to as a manipulation attack against RAG. Previous studies have explored various forms of adversarial manipulation attacks, such as adversarial attack on the retriever (liu_order-disorder_2023; liu2023black; wu2023prada), prompt injection attack (cai2022badprompt; liu2023prompt; jain2023baseline), jailbreak attack for LLM (deng2023jailbreaker; li2023multi; zhao2024weak), and poisoning attack targeting the retrieval corpus in RAG (zou2024poisonedrag; xue2024badrag).
Figure 1. As demonstrated by an example from our pilot study (see left box), we observed that ChatGPT can be induced through user prompts to fully replicate the context segments retrieved in its responses. Motivated by this observation, we can imitate and transparentize the retriever of the black-box RAG model by enumerating critical queries and candidates. The right part illustrates the opinion manipulation attack against black-box RAG-like systems. When queries on controversial topics are submitted to a benign black-box RAG, it typically retrieves documents reflecting diverse viewpoints and generates neutral, objective responses. Conversely, by subtly modifying a minimal number of documents using triggers generated via a surrogate retrieval model, which imitates the black-box retriever, the manipulated RAG exclusively retrieves documents endorsing a specific opinion. This constrains the LLM to generate biased responses and influences users’ opinions.
@@
However, previous studies (zou2024poisonedrag; xue2024badrag; cho2024typos) have primarily focused on attacks against RAG systems, which lack feasibility and have fundamental limitations. They mostly address white-box scenario or employ heuristic-based black-box attacks, without thoroughly analyzing the vulnerabilities within the retrieval stage. Furthermore, earlier research has mainly concentrated on factoid question answering tasks, such as questions like “Who is the CEO of OpenAI?” Those factual queries are less impactful, as RAG systems combined with fact-checking and value alignment mechanisms in LLMs are capable of rectifying these errors (xiang2024certifiably; tu2025rbft).
@@
Differently, the scenario explored in this paper primarily addresses controversial and opinion-based inquiries in RAG, such as “Should abortion be legal?” These questions are open-ended and require advanced levels of logical reasoning and summarization capabilities from LLMs. Current research on controversial topics is limited. In scenarios involving controversial topics, opinion manipulation can create “Information Bubbles”, which subsequently lead to the homogenization of user opinions (zhang2023homogenizationdilemma). In such cases, individuals’ views become easily swayed by the stance of the information they are exposed to. Attacks that manipulate opinions on such questions could potentially inflict deeper harm (epstein_search_2015). Through carefully crafted input, attackers can influence the orientation of content generated by RAG models, thus jeopardizing users’ cognitive processes and decision-making abilities.
@@
This paper primarily investigates adversarial opinion manipulation targeting the retriever in black-box RAG-like systems, which aligns with realistic and practical scenarios. The threat model presented here can be characterized as follows: the adversary can only query the RAG-like system and does not have access to the complete knowledge base or corpus, the retriever, or the parameters of the RAG. The attacker is only capable of injecting limited adversarially modified candidate texts into the corpus, while the retriever and the LLM remain black-boxed, intact, and unmodifiable.
@@
To address aforementioned challenges, in this paper, we propose FlippedRAG, a black-box attack method, to explore the reliability of RAG in controversial topics and investigate its impact on user cognition, as shown in Figure 1. By employing black-box retriever imitation (liu_order-disorder_2023) for opinion manipulation, we reveal specific vulnerabilities in RAG-like systems. Our motivation is based on the observation in Figure 1. When we queried ChatGPT (as of March 2025, with web search enabled, functioning as a typical RAG-like application) on a specific topic and explicitly instructed it to output the original content of its retrieved references, it faithfully reproduced the exact text segments from the retrieved candidates. The reproduced outputs precisely matched the corresponding content on the web pages linked in the reference citations and reflect the functionality of relevance matching. It allow us to designed a retriever imitation method for black-box RAG to address the limitations of white-box attacks or heuristic-based attacks against such systems. Specifically, we utilize carefully designed instructions to obtain the retrieval results within the RAG model. Subsequently, based on the constructed (query, candidate) pairs from retrieval list, we train a surrogate model using the extracted retrieval data (liu_order-disorder_2023; wu2023prada) to approximate and transparentize the relevance preferences of the retriever in the black-box RAG model.
@@
Based on this surrogate model, we develop an attack strategy aimed at manipulating the opinions of candidate documents. By attacking this surrogate model, we generate adversarial opinion manipulation triggers, which are then transferred to the target RAG corpus, as shown in the right part of Figure 1. We conduct experiments on opinion datasets encompassing multiple topics to validate the effectiveness and scope of the attack strategy without relying on internal knowledge of the RAG model. The experimental results reveal that the proposed attack strategy can significantly alter the opinion polarity of the content generated by RAG. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack. This not only highlights the vulnerability of the retrieval model and RAG model but, more importantly, underscores the potential negative impact on user cognition and decision-making. Such vulnerabilities can increase the risk of misleading users into accepting incorrect or biased information.
@@
(1) We propose FlippedRAG, a novel transfer-based opinion manipulation attack against black-box RAG-like systems. It realistically reflects practical threat scenarios, and to our knowledge, we are among the first to study adversarial opinion manipulation targeting open-ended and controversial topics in RAG-based models.
(2) We propose a new approach to construct high-quality pseudo-relevant contrastive pairs for training surrogate retrieval models in black-box RAG. It highlights risks regarding context data leakage in LLM applications. Extensive automated evaluations combined with user studies confirm that our method effectively manipulates the opinions generated by RAG models, ultimately influencing user perception toward specifically targeted stance.
@@ 2.1. Adversarial Retrieval Attack Strategy
The retrieval model is a core component in RAG. Many existing research evaluates its robustness and security by inducing faulty behaviors through adversarial attacks. The adversarial retrieval attack strategy starts with manipulation at the word level. Under white-box setting, Ebrahimi et al. (ebrahimi2017hotflip) utilize an atomic flip operation, which swaps one token for another, to generate adversarial text, named HotFlip. HotFlip eliminates reliance on rules, but the adversarial trigger it generates usually has incomplete semantics and insufficient grammar fluency. While it can deceive the target model, it cannot evade perplexity-based defenses. Wu et al. (wu2023prada) also proposed a word substitution ranking attack method called PRADA. To enhance the readability and effectiveness of the trigger, Song et al. (song2020adversarial) propose an adversarial method under white-box setting, named Collision, which uses gradient optimization and beam search to produce the adversarial trigger. Collision method further imposes a soft constraint on collision generation by integrating a language model, reducing the perplexity of the collision. Furthermore, Liu et al. (liu_order-disorder_2023) propose the Pairwise Anchor-based Trigger (PAT) generation under black-box setting. By adding the fluency and semantic consistency constraints, it generates adversarial triggers by optimizing the pairwise loss of top candidates and target candidates with triggers.
@@ 2.2. Reliability of RAG
Some studies have investigated the adversarial poisoning attacks on RAG models, including PoisonedRAG(zou2024poisonedrag), BadRAG(xue2024badrag), Phantom(chaudhari2024phantom), and GARAG(cho2024typos). The distinctive characteristics of these attack methodologies are summarized in the Table 1. “LLM Capability Utilization” refers to the motivation of leveraging LLMs’ capabilities to induce malicious outputs. “Depth” denotes the cognitive penetration depth of RAG attacks. Although both BadRAG(xue2024badrag) and Phantom(chaudhari2024phantom) target user perception, their manipulations remain confined to sentiment level without achieving deeper ideological persuasion.
As RAG is designed to overcome the hallucination problem in LLMs and enhance their generative capabilities, the reliability of content generated by RAG also becomes a major concern. Zhang et al. (zhang2024human) sought to identify the weaknesses of RAG by analyzing critical components in order to facilitate the injection of the attack sequence and crafting the malicious documents with a gradient-guided token mutation technique. Xiang et al. (xiang2024certifiably) proposed an isolate-then-aggregate strategy, named RobustRAG, which first obtains responses from each passage of the LLM in isolation and then securely aggregates these isolated responses, to construct the first defense framework against retrieval corruption attacks. These studies are based on white-box scenarios and primarily focus on the robustness of RAG against corrupted and toxic content.
Table 1. Comparison of existing RAG attacks. DoS denotes denial-of-service and "-" denotes not applicable.
@@
PoisonedRAG (zou2024poisonedrag) Closed-ended, Factoid White-box/Black-box Logical reasoning
BadRAG (xue2024badrag) Open-ended, Factoid, DoS White-box Logical reasoning Sentiment
Phantom (chaudhari2024phantom) Open-ended, Factoid, DoS White-box Instruction-following Sentiment
GARAG (cho2024typos) Closed-ended, Factoid Gray-box Noise resistance
FlippedRAG Controversial, Open-ended Black-box
Induction, Logical reasoning
Perspectives
In this paper, we propose the use of adversarial retrieval attack to manipulate the retrieval process in black-box RAG, ensuring that corrupted documents with a specific opinion are ranked as highly as possible, thereby inducing the LLM responses to align with that opinion on controversial topics.
@@ 3.1. Overview
Our objective is to manipulate the opinions expressed in the responses generated by black-box RAG models on controversial topics. We mainly focus on the retrieval component, where manipulated ranking outcomes propagate to bias the LLM’s output generation. Zhang et al. (zhang2024human) attempted to poison context documents to mislead the LLM into generating incorrect content. However, this approach necessitates extensive internal details of the LLM application, rendering it less feasible in real-world scenarios. In the black-box RAG context, the attacker lacks access to internal information of the RAG, including model architecture and scoring functions, and can only interact with the inputs and outputs of the RAG. Specifically, the attacker can only use the interface of the RAG and not directly access the retriever. Since the inputs to LLMs are comprised of the user query and context documents, and the user query is immutable, our attack strategy focuses on modifying candidate documents within the corpus. Although the attacker does not have access to the entire corpus, they can insert adversarially modified candidate texts into it, as many RAG-like applications source information from the Internet, where the content is publicly accessible and editable.
The basic framework of RAG comprises two components: the retriever and the LLM. These modules are serially connected, with the retriever sourcing context information from a knowledge base, upon which the LLM then performs the generation task. In black-box scenarios, attackers cannot modify the system prompts of the generative LLM, making it challenging to directly influence the generation results by exploiting any reliability flaws within the LLM itself. Therefore, we focus on exploiting the reliability flaws of the retriever to manipulate the retrieval ranking results.
We employ specific instructions to induce RAG to reveal the context information it references. With this information, we train a surrogate model that approximates the black-box retriever, effectively transparentizing it into a white-box model. We then employ this white-box model to generate adversarial triggers. By adding these adversarial triggers to candidate documents that reflect the target opinion, we enhance their relevance to the user query, increasing the likelihood that they will be included in the context passed to the generative LLM. Leveraging the strong capability of LLM for context understanding, summarization and instructions-following, we guide it to generate responses that align with the target opinion. An overview of this process is shown in Figure 2.
Figure 2. The overview of FlippedRAG for manipulating the opinions of RAG-generated content under black-box setting.: